Write the equation of a line with slope -4/5 and a y intercept of 4.
Answer by ewatrrr(24785)


Hi,
Using the standard slope-intercept form for the equation of a line
y = mx + b where m is the slope and b the y-intercept
equation of a line with slope -4/5 and a y intercept of 4.
y =(-4/5)x + 4
standard form of that line would be
(4/5)x + y = 4
multiplying each term on side of the equation by 5
4x +5y = 20